OpenStudy (alexandervonhumboldt2):

I offer help with Al 1 and 2 or lower and Geometry too

OpenStudy (anonymous):

how many solutions does the system of equations have? 15y = -3x + 9 y = -1/5x + 9/15 a. 1 b. 2 c. none d. infinitely many

OpenStudy (anonymous):

I have a question too when yall are done :)

OpenStudy (anonymous):

I would say to because there are infinitely many because the lines are equal

OpenStudy (alexandervonhumboldt2):

ok post it

OpenStudy (anonymous):

y=-3/15x+9/15=-1/5x+9/15. this one equals the second one, so the system has infinitely many solutions

OpenStudy (anonymous):

what 's your question @lenalovey

OpenStudy (alexandervonhumboldt2):

listen sister i will help go away